OpenStudy (anonymous):

Find the following values for the given functions. f(x) = x + 5, g(x) = x^2 (f+g)(5)=..?

OpenStudy (driftracer305):

my brother..... uchiha

OpenStudy (anonymous):

well, (f+g)(x) = f(x) + g(x)

OpenStudy (anonymous):

now can u do the rest?

OpenStudy (anonymous):

hwahhaha.. lol @driftracer305 .. hwaha.. @PFEH.1999 so i will first substitute the 5 to f(x) and g(x) before i perform the solution..?

OpenStudy (driftracer305):

k first always plug in given value like 5 in the last letter of equation like you have g....then plug the result as a value of x for f(x)

OpenStudy (anonymous):

ok, first let's see what's (f+g)(x) f(x) + g(x) = x^2 + x + 5

OpenStudy (anonymous):

and then you can find (f+g)(x) ;)

OpenStudy (anonymous):

oh.. so i will make it as x^2 + x+ 5 before doing the substituting the 5..?

OpenStudy (anonymous):

sorry wrong grammar.. hahah

OpenStudy (anonymous):

i told you to do this because it would help you to understand but you can first plug 5 too ;)

OpenStudy (driftracer305):

i always do it as g(5) = 5^2 =25........... now you have f(25) = 25 + 5 =30

OpenStudy (anonymous):

you can say what is f(5) and g(5) and then find f(5) +g(5) ;)

OpenStudy (zzr0ck3r):

addition is linear operator so \((f+g)(x)=f(x)+g(x)\)

OpenStudy (anonymous):

yes @zzr0ck3r but i told this ;)

OpenStudy (anonymous):

ahh.. yeah.. but will it be the same if i add it first before substituting or substitute first before add.?

OpenStudy (anonymous):

yes ;)

OpenStudy (zzr0ck3r):

ie plug in after = plug in before :)

OpenStudy (zzr0ck3r):

nope, same thing

OpenStudy (zzr0ck3r):

thats the cool part!

OpenStudy (driftracer305):

i dont knw what u jus said @russeldano ........ but like i said... plug and evaluate............ KEEP Calm and be an Uchiha....... lol

OpenStudy (anonymous):

@driftracer305 yeah.. i was thinking about that but i wasnt sure so i asked.. hahah

OpenStudy (zzr0ck3r):

@russeldano try it both ways to convince yourself

OpenStudy (driftracer305):

yea...... math has one answer but thousand ways to get it..... so do what u fell good with

OpenStudy (zzr0ck3r):

nah, once you feel good, try it different ways. but do it the one way on the test:)

OpenStudy (anonymous):

@zzr0ck3r woah.. haha.. its the same.. awesome.. hwahh.. i get it now.. hwahha.. @driftracer305 yeah.. thats why i like math.. but sometimes i forgot someways to do the equation.. hwhaha.. tnx @zzr0ck3r , @PFEH.1999 , @driftracer305 .. haha..

OpenStudy (anonymous):

yes,after solving a problem . try other ways . it would help u to understand questions better and solve harder one ;) you're welcome :)

OpenStudy (anonymous):

last question.. is this correct..? (fg)(5)= f(5)*g(5)..? haha...

OpenStudy (driftracer305):

lol...... u could have just used your sharingan @russeldano ...... if you know wat i mean.. ;)

OpenStudy (anonymous):

tes;)

OpenStudy (anonymous):

hwahaha.. @driftracer305 i guess i know what u mean there.. hwahha

OpenStudy (driftracer305):

ohh yea.. it is right

OpenStudy (anonymous):

(f+g)(x) = f(x) + g(x) (fg)(x) = f(x) . g(x) (f/g)(x) = f(x) / g(x)

OpenStudy (anonymous):

and (f-g)(x) = f(x) - g(x) too ;)

OpenStudy (anonymous):

ahh.. got it know.. but can i ask again a question.. will i perform this or just substitute..? Find f (a), f (a + h), and the difference quotient for the function given below, where h not equal to 0. f (x) = 4x2 + 7 f(a+h)-f(a)/h =..?

OpenStudy (anonymous):

well,ok.let's start with this one: f(a) = a^2 + a and then f(a+h) = ? let's imagine a+h = y so we should find f(y) f(y) = y^2 + y and then y = a+h so f(a+h) = (a+h)^2 + (a+h) ;) got it?

OpenStudy (anonymous):

didnt get the part where f(a) = a^2 + a..? where did that came from..?

OpenStudy (anonymous):

it is only an example ;)

OpenStudy (anonymous):

i wanted you to know this and then solve that question

OpenStudy (anonymous):

ooopss sorry f (x) = 4x2 + 7 it was suppose to be f (x) = 4x^2 + 7 haha ahh.. ok got it..

OpenStudy (anonymous):

and now can you solve it yourself ?

OpenStudy (anonymous):

ok ok got it.. then.. haha

OpenStudy (anonymous):

nice :)

OpenStudy (anonymous):

wait i will show you what is my answer ok..?

OpenStudy (anonymous):

(4(a+h)^2 + 7)-(4a^2 +7) / h..?? im stuck at here.. dont know what is the next move..? haha..

OpenStudy (anonymous):

it's correct :)

OpenStudy (anonymous):

thats the final answer..? no need to do any simplification..?

OpenStudy (anonymous):

no,the question wants us f(a+h)-f(a)/h and u found it ;) any other questions?

OpenStudy (anonymous):

ahh.. ok.. no need to simplify.. just follow the direction as is.. haha.. tnx.. you've really helped me.. hahah

OpenStudy (anonymous):

you're welcome ;) functions are really interesting